OpenStudy (anonymous):

Find the product of (x2 + 4x ��������������������������� 3)(2x2 + x + 6) hey @birvin73 could you help me? will medal

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Multiply all the terms in the first parenthesis to all the terms in the 2nd parenthesis. x^2 * 2x^2 = ? x^2 * x = ? x^2 * 6 = ? 4x * 2x^2 = ? 4x * x = ? 4x * 6 = ? -3 * 2x^2 = ? -3 * x = ? -3 * 6 = ?

OpenStudy (anonymous):

2x^3 x^3 6x^2 8x^3 4x^2 24x -6x^2 -3x -12

OpenStudy (igreen):

Yep, all are correct except the first and last one.

OpenStudy (igreen):

x^2 * 2x^2 = ? -3 * 6 = ?

OpenStudy (anonymous):

oh -18 2x^4

OpenStudy (igreen):

Yep, now here's the whole thing: 2x^4 + x^3 + 6x^2 + 8x^3 + 4x^2 + 24x - 6x^2 - 3x - 12 Combine like terms, can you do that?

OpenStudy (anonymous):

so 2x^4+9x^3+4x^2+21x-12

OpenStudy (igreen):

Yep, you got it!

OpenStudy (anonymous):

thank u

OpenStudy (igreen):

No problem.